how many 60lb bags of concrete do i need to make a wall 15ftlg x15in high x 3in thick
Found 2 solutions by ikleyn, jim_thompson5910:
Answer by ikleyn(52832)

.
Do you know what concrete is?

Read this Wikipedia article.


Answer by jim_thompson5910(35256)

1 ft = 12 inches
15*1 ft = 15*12 inches
15 ft = 180 inches

The wall of "15 ft x 15 in x 3 in" is the same dimensions as the wall of "180 in x 15 in x 3 in"

Use the volume formula
V = L*W*H
V = 180*15*3
V = 8100

The volume of the wall is 8100 cubic inches

Here's where I can't keep going. I don't know how many cubic inches one bag can make. You'll have to look at each bag to determine this. That kind of information should be given.

Let's say for example that each bag can make 100 cubic inches of concrete. That would mean you need 8100/100 = 81 bags. Replace "100" with whatever the bag states. If you get a decimal result, round up (not down) to the nearest whole number.